IQs are calibrated so that they have a mean of 100 and a standard deviation of 15.
a) Find the probability that a randomly selected person has an IQ between 106 and 116. Be sure to draw a sketch.
b) A person with an IQ at or below the 5th percentile is considered “developmentally disabled”. What IQ corresponds to the cutoff for this designation?

Answer by Theo(13342)

the easiest way to figure this out is to use the following online statistical z-score calculator.

http://davidmlane.com/hyperstat/z_table.html

the probability that a randomly selected person has an IQ between 106 and 116 is shown below, using this calculator:

$$$

the answer shown is .2015

the IQ at or below the 5th percentile is shown below:

$$$

the answer shown is 75.322

this calculator allows you to analyze the distribution without having to translate to z-scores.

if you did have to translate to z-scores, you would have done the following, using the formula indicated.

z = (x-m)/s

z is the z-score
x is the raw score
m is the mean
s is the standared deviation.

in the first part of your problem, you would calculate two z-scores.

z1 is the low z-score of the range.
z2 is the high z-score of the range.

you would have gotten:

z1 = (106-100)/15 = .4

z2= (116-100)/15 = 1.067

you would then look up the area on the distribution curve between these z-scores.

the same calculator, using z-scores rather than raw scores, would assist you to find the answer.

that would be shown below:

$$$


the4 answer is the same as before, i.e. = .2015

you would also use the same calculator using a mean of 0 and a standard deviation of 1 to finf the z-scpore associated with the fifth percentile of IQs.

that z-score would be -1.645 as shown below:

$$$


you would then find the raw score by using the z-score formula as below:

z = (x-m)/s

-1.645 = (x - 100) / 15

solve for x to get x = -1.645 * 15 + 100 which is equal to 75.325

any small difference is more then likely due to rounding.

the z-score is a normalized score.

the mean becomes 0 and the standard deviation becomes 1.

a z-score of 0 is the same as the mean of the raw score.
a standard deviation of 1 is the same as the standard deviation of the raw score.

this means that a z-core is the same as a raw score of 100 in your problem, and a z-score of 1 is the same as a raw score of 115 in your problem, and a z-score of -1 is the same as a z-score of 85 in your problem.